# Revisiting Silhouette: From Micro to Macro Aggregation

### 11 Jan 2024 v1

@@ Abstract
Silhouette coefficient is an established internal clustering evaluation measure that produces a score per data point, assessing the quality of its clustering assignment. To assess the quality of the clustering of the whole dataset, the scores of all the points in the dataset are typically averaged into a single value, a strategy which we call as micro-averaging. As we illustrate in this work, by using a synthetic example, this micro-averaging strategy is sensitive both to cluster imbalance and outliers (background noise). To address these issues, we propose an alternative aggregation strategy, which first averages the silhouette scores at a cluster level and then (macro) averages the scores across the clusters. Based on the same synthetic example, we show that the proposed macro-averaged silhouette score is robust to cluster imbalance and background noise. We have conducted an experimental study showing that our macro-averaged variant provides better estimates of the ground truth number of clusters on several cases compared to the typical micro-averaged score.
@@ 1 Introduction
In this study, we advocate averaging the silhouette values across clusters (macro-averaging) and not across the individual data points (micro-averaging) to obtain a single score for evaluating the clustering assignments of the entire dataset. Although the latter is the default aggregation strategy, we show that by assigning equal weight to each cluster, the former is more effective in imbalanced or noisy cases.
To substantiate our claim, we have conducted a series of experiments on synthetic imbalanced and noisy clustering problems, by estimating the number of clusters in eight real-world datasets. Our results show that the proposed macro-averaged silhouette coefficient is more robust concerning noise, imbalance, and sample size.
@@ 2 The Silhouette Coefficient
A large $b(x_{i})$ value indicates that $x_{i}$ significantly differs from the points of the closest cluster. In Figure 1, the closest cluster (which minimizes $b(x_{i})$) is $C_{J}$. Considering both $a(x_{i})$ and $b(x_{i})$, the silhouette score of $x_{i}$ is defined as:
@@ 3 Aggregating Silhouette
Since the silhouette coefficient provides a score that grades the cluster assignment of a data point, in order to obtain a single score for the whole partition of dataset $X$, the individual silhouette scores are averaged at the data point level. We will refer to the aggregation strategy at the data point level as micro-averaging, defined as follows:
@@
The scikit-learn implementation of the silhouette coefficient employs micro-averaging by default, and several studies have used it to estimate the number of clusters. However, as we show in this work, this aggregation strategy has two weaknesses. First, it is not effective in the case of imbalanced clusters. Second, it is sensitive to the presence of background noise, i.e., data points lying in the space between the clusters. Both of these weaknesses can be addressed by using the macro-averaged silhouette coefficient, which we define next.
3.1 Macro-Averaged Silhouette
@@
When clusters are perfectly balanced, the sample mean (i.e., micro-averaging) is a reasonable aggregation strategy. This assumption, however, of clusters being perfectly balanced, cannot be guaranteed in the real world, where clusters are often imbalanced (see, for example, Table 1). In this case, micro-averaging is not effective as discussed below in more detail. To address this issue, we propose a cluster-level aggregation strategy of the individual silhouette scores, which we will refer to as macro-averaging. Specifically, a silhouette score $S_{c}$ is computed for each cluster $C_{i}$ as follows:
@@
Hence, for each cluster $C_{i}$ we compute a single silhouette value $S_{c_{i}}$. This value can be interpreted as a score that measures how compact and well separated a cluster given a clustering solution. Consequently, we end up with a set of $K$ silhouette values {$S_{C_{1}},\ldots,S_{C_{K}}$}, where $K$ is the number of clusters in the solution. To evaluate the clustering solution as a whole, we use the mean to aggregate these $K$ scores into a single value. Although the mean is a straightforward option, we note that other statistics can also be used for this aggregation (e.g., min, max, median). More formally, our proposed silhouette variant is defined as follows:
@@
By equally weighing the clusters, we show next how this macro-aggregation strategy overcomes the cluster imbalance issue.
Dataset Type Description ${N}$ ${d}$ ${k}$ $r$ Source
pendigits Time-series Handwritten digits 10,992 16 10 0.92 [17]
mnist (test) Image Handwritten digits 10,000 28$\times$28 10 0.79 [18]
optdigits Image Handwritten digits 5,620 8$\times$8 10 0.97 [17]
mice Tabular Expression levels of proteins 1,080 77 8 0.70 [17]
olivetti faces Image Human Faces 400 64$\times$64 40 1.00 at&t
tcga Tabular Cancer gene expression profiles 801 20,531 5 0.26 [17]
glass Tabular Types of glass 214 9 6 0.12 [17]
wine Tabular Chemical analysis of wines 178 13 3 0.68 [17]
An imbalanced clustering example
Consider the synthetic dataset shown in Figure 1(a), which consists of $12$ Gaussian clusters with $100$ data points each. The clusters are of different variance, ranging from low to high values. Figure 1(b) presents the perfect clustering, where points in the same cluster are coloured similarly. The typical (micro-averaged) silhouette score $S_{micro}$ of this perfect clustering is $0.74$. Figure 1(c) shows the same data points as in Figure 1(b), but a random cluster assignment is hypothesised for all data points, except for those belonging to the compact low-variance cluster (nucleus) lying at the center of the dataset. For this suboptimal clustering, $S_{micro}$ drops to $-0.20$. Then, we disturb the cluster balance by adding points to the nucleus cluster, using a very small variance to preserve its dense structure, and we evaluated how the increase in cluster imbalance affects $S_{micro}$ score. Figure 3 shows that the score increases monotonically with the nucleus size and it even exceeds the score of the correct partition for size above $5,000$ data points. $S_{macro}$, on the other hand, has a value of $-0.20$ at all cases of this random cluster assignment and a value of 0.74 for the perfect clustering.
Figure 3: Micro- and macro-averaged silhouette (vertically) as a function of the size of the nucleus cluster (Figure 1(c)). The score for the optimal clustering (Figure 1(b)) is shown in yellow.
Cluster-balanced sampling for silhouette computation
@@
To solve this problem, following the intuition of our proposed macro-averaged silhouette score, we propose creating a subsample (of size $L$) for computing the macro-averaged silhouette score by uniformly selecting a subset of $L/K$ points from each cluster $C_{i}$, where $i=1,\ldots,K$. In this way, we ensure that all the clusters contribute a sufficient number of data points to the subsample, and a robust macro-averaged silhouette score computation.
@@
Estimating the true number of clusters
One of the typical uses of silhouette is to estimate the number of clusters. This is achieved by solving the clustering problem for several $K$ values and selecting the solution with the maximum silhouette score.
@@
Figure 5 shows the silhouette score of the solutions for varying $k$, when it is macro- (green line) and micro-averaged (red line). A star indicates the maximum per line. It is clear the the proposed macro-variant always finds the true number of clusters (yellow dashed vertical line), despite the heavy cluster imbalance. In contrast, the default micro-averaged scores are saturated near $0.95$ and a plateau for $k\in[6,15]$) with no clear peak being obvious, which leads to a wrong estimation for the true number of clusters.
Sensitivity to background noise
Besides its sensitivity on cluster imbalance, already shown in the previous paragraphs, we present here another weakness of the typical(micro-averaged) silhouette score which is related to the existence of background noise, i.e., points lie between clusters or far from the clusters. Figure 5(a) depicts four well-separated Gaussian clusters, each comprising $200$ data points. We introduced background random noise to this dataset, by sampling $n$ points from a uniform distribution $\epsilon_{n}\sim\mathcal{U}(a,b)$, with $a,b$ chosen so that the samples are spread uniformly in the data space, as shown in Figure 5(d). If $n$ noisy points are added, we define the relative level of background noise that is added to a dataset $X$ as $\frac{n}{n+|X|}\times 100\%$, As an example, Figure 5(b) (5(c)) presents the original dataset with 25% (50%) of noisy points added. We generated several datasets (i.e., versions of the original dataset) by increasing the noise level. To each such dataset we applied global $k$-means++, with $k$ ranging from 1 to 30, and we computed the micro- and macro-averaged silhouette scores per $k$. The $k$ with the highest score was considered as the estimated correct number of clusters per dataset. Figure 5(d) presents the number of clusters estimation with respect to the noise level. It is impressive that the macro score always provides the correct number of clusters. In contrast, the estimations using the micro score are severely affected as the noise level increases.
@@
Figure 7: Micro- (red) and the macro-averaged (green) silhouette score per dataset, evaluating the partitions of $k$-means for varying $k$ values.
@@ 4.2 Experimental results
As with the synthetic dataset we have used the global $k$-means++ algorithm [20] to cluster each dataset for increasing values of $k$. Then, we measured the silhouette score for each solution ($k$), by employing both (the default) micro- and (the proposed) macro-averaging. The best clustering solution with respect to each score is selected for each dataset. All the datasets comprise ground truth ($k$ in Table 1), hence it is reasonable to expect that the maximum silhouette score would be observed at that number of clusters, disregarding the aggregation strategy.
@@
Results
Figure 7 depicts the micro- and the macro-averaged silhouette score per dataset. When we used the macro-averaging strategy (green line), the optimal score (starred) coincides with the ground truth (the yellow dashed line) in five out of eight datasets (pendigits, olivetti faces, glass, mice, wine), covering all the different data types (time-series, image, tabular data). When using micro-averaging (red line), the ground truth $k$ was correctly estimated in three datasets (mice, wine, optdigits). These three datasets cover two data types (image, tabular) and for the latter ones the ground truth was also correctly estimated when macro-averaging (overlapping stars). In one tabular dataset (tcga), both strategies failed to capture the ground truth. In two datasets (optdigits, tcga), the maximum macro-averaged silhouette was observed for a $k$ other than the ones defined in the ground truth. In optdigits, where the ground truth is defined as one cluster per digit ($k=10$), the best macro-average score is established when the digits 3 and 9 are clustered together. A similar observation can be drawn by looking at the red curve, where the best micro-averaged silhouette score is in a plateau covering two solutions, i.e., $k\in\{9,10\}$. A similar red plateau is observed for the tcga dataset, where the best solution could be any from a range, i.e., $5\leq k\leq 8$. On the contrary, macro-averaging suggests a solution ($k=3$) that falls out of that range. Interestingly, however, the second highest peak is at the true number of clusters (i.e., 5).
There are also cases where the score is not effective. Such a case can be observed in the mnist dataset where both aggregation strategies prefer the solution with $K=2$, while the ground-truth number of clusters is $K=10$. We note, however, that even the ground truth solution yields very low silhouettes scores ($S_{micro}\approx 0.049$, $S_{macro}\approx 0.045$). Although similar to optdigits in nature (i.e., images of nine handwritten digits), the complexity of the mnist data is higher (i.e., more dimensions and complicated digit patterns), yielding aggregated silhouette scores close to zero, thus both strategies fail in estimating the correct number of clusters.
@@ 5 Conclusions
In this work we have illustrated two key weaknesses of the widely-used silhouette score for evaluating clustering solutions. Based on a synthetic example, we demonstrated the sensitivity of the default micro-averaging strategy to cluster imbalance and background noise. To address these issues, we proposed an alternative aggregation strategy, which measures the silhouette coefficient per cluster, and then returns their macro-average. As empirically shown, the proposed macro-averaged silhouette score is robust to imbalance and background noise, even when subsamples of the dataset are considered. Furthermore, we have applied clustering on eight well-known real-world datasets, for various values of $K$, and we assessed the two silhouette aggregation strategies regarding their ability to yield a maximum silhouette score for the ground truth number of clusters. The presented experimental results show that the proposed macro-averaged silhouette score leads to better estimations for several cases, clearly indicating that it constitutes a serious alternative to the typical silhouette coefficient.
Code availability. The code is publicly available at: https://github.com/gvardakas/macro-silhouette.

### 15 Jan 2024 v2

# Silhouette Aggregation: From Micro to Macro

@@ Abstract
Silhouette coefficient is an established internal clustering evaluation measure that produces a score per data point, assessing the quality of its clustering assignment. To assess the quality of the clustering of the whole dataset, the scores of all the points in the dataset are either (micro) averaged into a single value or averaged at the cluster level and then (macro) averaged. As we illustrate in this work, by using a synthetic example, the micro-averaging strategy is sensitive both to cluster imbalance and outliers (background noise) while macro-averaging is far more robust to both. Furthermore, the latter allows cluster-balanced sampling which yields robust computation of the silhouette score. By conducting an experimental study on eight real-world datasets, estimating the ground truth number of clusters, we show that both coefficients, micro and macro, should be considered.
@@ 1 Introduction
In this study, we advocate averaging the silhouette values both across clusters (macro-averaging) and across data points (micro-averaging) to obtain a single score for evaluating the clustering assignments of the entire dataset. The micro-averaged silhouette index is the originally proposed aggregation strategy [1] and the default implementation in scikit-learn, which have made it a very popular approach, e.g., to define the number of clusters. In this work, we show that by assigning equal weight to each cluster, the former is more effective in imbalanced or noisy cases.
To substantiate our claim, we have conducted a series of experiments on synthetic imbalanced and noisy clustering problems, and by estimating the number of clusters in eight real-world datasets. Our results show that the macro-averaged silhouette coefficient is more robust concerning noise, imbalance, and sample size.
@@ 2 The Silhouette Coefficient
A large $b(x_{i})$ value indicates that $x_{i}$ significantly differs from the points of the closest cluster. In Figure 1, the closest cluster (which minimises $b(x_{i})$) is $C_{J}$. Considering both $a(x_{i})$ and $b(x_{i})$, the silhouette score of $x_{i}$ is defined as:
@@ 3 Aggregating Silhouette
Since the silhouette coefficient provides a score that grades the cluster assignment of a data point, in order to obtain a single score for the whole partition of dataset $X$, the individual silhouette scores have so far either (micro) averaged at the data point level or macro-averaged at the cluster level.
3.1 The Micro-averaged Silhouette Index
Micro-averaging silhouette at the point level is defined as follows:
@@ 3.1 The Micro-averaged Silhouette Index
This is the proposed averaging strategy by the study that introduced silhouette [1] and it is adopted by the widely-used library of scikit-learn, making it the typical approach employed in literature [17]. However, as we show in this work, this aggregation strategy has two weaknesses. First, it is not effective in the case of imbalanced clusters. Second, it is sensitive to the presence of background noise, i.e., data points lying in the space between the clusters.
3.2 The Macro-averaged Silhouette Index
@@
When clusters are perfectly balanced, the sample mean (i.e., micro-averaging) is a reasonable aggregation strategy. This assumption, however, of clusters being perfectly balanced, cannot be guaranteed in the real world, where clusters are often imbalanced (see, for example, Table 1). In this case, micro-averaging is not effective, as will be discussed next in more detail. By contrast, the cluster-level (macro) aggregation strategy is more robust. In this strategy, a silhouette score $S_{c}$ is computed for each cluster $C_{i}$ as follows [^1] We note that the R implementation of the silhouette score employs macro-averaging. To the best of our knowledge, very few studies in literature deviate from the typical approach and follow this paradigm [18].:
@@
Hence, for each cluster $C_{i}$ a single silhouette value $S_{c_{i}}$ is computed, known as cluster mean silhouette. This value can be interpreted as a score that measures how compact and well separated a cluster given a clustering solution. Consequently, we end up with a set of $K$ silhouette values {$S_{C_{1}},\ldots,S_{C_{K}}$}, where $K$ is the number of clusters in the solution. To evaluate the clustering solution as a whole, the mean aggregates these $K$ scores into a single value. Although the mean is a straightforward option, we note that other statistics can also be used for this aggregation (e.g., min, max, median). More formally, this macro-averaged silhouette variant (i.e., equally weighing the clusters) is defined as follows:
@@
3.3 An imbalanced clustering example
Consider the synthetic dataset shown in Figure 1(a), which consists of $12$ Gaussian clusters with $100$ data points each. The clusters are of different variance, ranging from low to high values. Figure 1(b) presents the perfect clustering, where points in the same cluster are coloured similarly. The micro-averaged silhouette score $S_{micro}$ of this perfect clustering is $0.74$. Figure 1(c) shows the same data points as in Figure 1(b), but a random cluster assignment is hypothesised for all data points, except for those belonging to the compact low-variance cluster (nucleus) lying at the center of the dataset. For this suboptimal clustering, $S_{micro}$ drops to $-0.20$. Then, we disturb the cluster balance by adding points to the nucleus cluster, using a very small variance to preserve its dense structure, and we evaluated how the increase in cluster imbalance affects the $S_{micro}$ score. Figure 3 shows that the score increases monotonically with the nucleus size and it even exceeds the score of the correct partition for size above $5,000$ data points. $S_{macro}$, on the other hand, has a value of $-0.20$ at all cases of this random cluster assignment and a value of 0.74 for the perfect clustering.
Figure 3: Micro- and macro-averaged silhouette (vertically) as a function of the size of the nucleus cluster (Figure 1(c)). The score for the optimal clustering (Figure 1(b)) is shown in yellow. In black is shown the highest possible silhouette score.
3.4 Cluster-balanced sampling for silhouette computation
@@
To solve this problem, following the intuition of the macro-averaged silhouette score, we propose creating a subsample of size $L$ for computing the macro-averaged silhouette score by uniformly selecting a subset of $L/K$ points from each cluster $C_{i}$, where $i=1,\ldots,K$. In this way, we ensure that all the clusters contribute a sufficient number of data points to the subsample, and a robust macro-averaged silhouette score computation.
@@
3.5 Estimating the true number of clusters
One of the typical uses of silhouette is to estimate the number of clusters. This is achieved by solving the clustering problem for several $K$ values and selecting the solution with the maximum silhouette score [20].
@@ 3.5 Estimating the true number of clusters
Figure 5 shows the silhouette score of the solutions for varying $k$, when it is macro- (green line) and micro-averaged (red line). A star indicates the maximum per line. The macro-averaged score always finds the true number of clusters (yellow dashed vertical line), despite the heavy cluster imbalance. In contrast, the default micro-averaged scores are saturated near $0.95$ and a plateau for $k\in[6,15]$. The lack of a clear peak leads to the wrong estimation for the true number of clusters.
3.6 Sensitivity to background noise
Besides its sensitivity on cluster imbalance, shown in the previous paragraphs, we present another weakness of the micro-averaged silhouette score that is related to the existence of background noise, i.e., points lie between clusters or far from the clusters.
Figure 5(a) depicts four well-separated Gaussian clusters, each comprising $200$ data points. We introduced background random noise to this dataset, by sampling $n$ points from a uniform distribution $\epsilon_{n}\sim\mathcal{U}(a,b)$, with $a,b$ chosen so that the samples are spread uniformly in the data space, as shown in Figure 5(d). If $n$ noisy points are added, we define the relative level of background noise that is added to a dataset $X$ as $\frac{n}{n+|X|}\times 100\%$, As an example, Figure 5(b) (5(c)) presents the original dataset with 25% (50%) of background noise.
We generated several datasets (i.e., versions of the original dataset) by increasing the noise level. To each such dataset we applied global $k$-means++, with $k$ ranging from 1 to 30, and we computed the micro- and macro-averaged silhouette scores per $k$. The $k$ with the highest score was considered as the estimated correct number of clusters per dataset. Figure 5(d) presents the number of clusters estimation with respect to the noise level, where it can be seen that the macro-averaged score always provides the correct number of clusters. In contrast, the estimations using the micro score are severely affected as the noise level increases.
@@ 3.6 Sensitivity to background noise
Figure 7: Micro- (red) and macro-averaged (green) silhouette score per dataset, evaluating the partitions of $k$-means for varying $k$ values.
@@ 4.1 Datasets
Dataset Type Description ${N}$ ${d}$ ${k}$ $r$ Source
pendigits Time-series Handwritten digits 10,992 16 10 0.92 [24]
mnist (test) Image Handwritten digits 10,000 28$\times$28 10 0.79 [25]
optdigits Image Handwritten digits 5,620 8$\times$8 10 0.97 [24]
mice Tabular Expression levels of proteins 1,080 77 8 0.70 [24]
olivetti faces Image Human Faces 400 64$\times$64 40 1.00 at&t
tcga Tabular Cancer gene expression profiles 801 20,531 5 0.26 [24]
glass Tabular Types of glass 214 9 6 0.12 [24]
wine Tabular Chemical analysis of wines 178 13 3 0.68 [24]
@@ 4.2 Experimental results
As with the synthetic dataset we have used the global $k$-means++ algorithm [21] to cluster each dataset for increasing values of $k$. Then, we measured the silhouette score for each solution ($k$), by employing both micro- and macro-averaging. The best clustering solution with respect to each score is selected for each dataset. All the datasets comprise ground truth ($k$ in Table 1), hence it is reasonable to expect that the maximum silhouette score would be observed at that number of clusters, disregarding the aggregation strategy.
@@
Correct number of clusters found
Figure 7 depicts the micro- and macro-averaged silhouette score per dataset. When we used the macro-averaging strategy (green line), the optimal score (starred) coincides with the ground truth (the yellow dashed line) in five out of eight datasets (pendigits, olivetti faces, glass, mice, wine), covering all the different data types (time-series, image, tabular data). When using micro-averaging (red line), the ground truth $k$ was correctly estimated in three datasets (mice, wine, optdigits). These three datasets cover two data types (image, tabular) and for the latter ones the ground truth was also correctly estimated when macro-averaging (overlapping stars).
Correct number of clusters not found
In one tabular dataset (tcga), both strategies failed to capture the ground truth. In two datasets (optdigits, tcga), the maximum macro-averaged silhouette was observed for a $k$ other than the ones defined in the ground truth. In optdigits, where the ground truth is defined as one cluster per digit ($k=10$), the best macro-average score is established when the digits 3 and 9 are clustered together. A similar observation can be drawn by looking at the red curve, where the best micro-averaged silhouette score is in a plateau covering two solutions, i.e., $k\in\{9,10\}$. A similar red plateau is observed for the tcga dataset, where the best solution could be any from a range, i.e., $5\leq k\leq 8$. On the contrary, macro-averaging suggests a solution ($k=3$) that falls out of that range. Interestingly, however, the second highest peak is at the true number of clusters (i.e., 5).
Silhouette ineffectiveness
There are also cases where the silhouette score is not effective with both aggregation strategies. Such a case can be observed in the mnist dataset where both aggregation strategies prefer the solution with $K=2$, while the ground-truth number of clusters is $K=10$. We note, however, that even the ground truth solution yields very low silhouettes scores ($S_{micro}\approx 0.049$, $S_{macro}\approx 0.045$). Although similar to optdigits in nature (i.e., images of nine handwritten digits), the complexity of the mnist data is higher (i.e., more dimensions and complicated digit patterns), yielding aggregated silhouette scores close to zero, thus both strategies fail in estimating the correct number of clusters.
@@ 5 Conclusions
In this work we revisited the widely-used silhouette coefficient, showing that the aggregation strategy should not be selected arbitrarily. Based on a synthetic example, we illustrated that the micro-averaged silhouette score for evaluating clustering solutions suffers from two key weaknesses due to its sensitivity to cluster imbalance and background noise. By contrast, as we show, the macro-averaged silhouette is robust to both these issues. Inspired by the macro-averaged strategy, we proposed a novel sampling strategy, which is considerably more robust compared to uniform sampling. By applying clustering on eight well-known real-world datasets, for various values of $K$, we assessed the two silhouette aggregation strategies regarding their ability to yield a maximum silhouette score for the ground truth number of clusters. Our experimental results show that the macro-averaged strategy leads to better estimations overall, but there is no clear winner, indicating that the use of both strategies when aggregating silhouette is the best option.